n1=1.00
n2=3.14
i=60 degrees
r=16 degrees
Air to wafflecheese
Whats the speed of light in wafflecheese
1
Expert's answer
2020-10-23T11:50:24-0400

The speed of ligh is expressed through given refractive indicex as following:


"\dfrac{c_{vacuum}}{c_{wafflecheese}} = n_2"

where "c_{vacuum} = 3\times 10^8 m/s" is the speed of light in vacuum and "n_2 = 3.14" is the refractive index of wafflecheese. Thus, obtain:

"c_{wafflecheese} = \dfrac{c_{vacuum}}{n_2}=\dfrac{3\times 10^8m/s}{3.14} \approx 0.96\times 10^8m/s"

Answer. "0.96\times 10^8m/s".